# A fluid–peridynamic structure model of deformation and damage of microchannels

### v1

@@ 4.3 Dispersion relation for the coupled problem
where $H_{0}$, $Q_{0}$, and $P_{0}$ denote the base state values, and $H^{\prime}$, $Q^{\prime}$, and $P^{\prime}$ represent the corresponding perturbations. Furthermore, to derive a linearized dispersion relation, we have assumed harmonic forms of the perturbations [52, 49, 50, 51] in Eqs. (22). Here, $H^{\prime}$, $P^{\prime}$, and $Q^{\prime}$ are understood as the real part of the complex expressions on the right-hand sides, but we leave the real part notation understood, as is common in the literature. In these expressions, $k$ is the (dimensionless) wavenumber, and $\omega$ is the (dimensionless) frequency.
@@
where the first row arises from the peridynamic beam equation, the second from the continuity equation, and the third row from the fluid’s axial momentum balance equation.
@@
The dispersion relation, Eq. (29), can be solved to determine the dimensionless phase velocity, $v_{p}=\Re{(\omega)}/\Re{(k)}$. Unlike the case of an uncoupled peridynamic beam or bar [52, 50, 51], where the dispersion relation depends solely on the wavenumber, the present model exhibits a more complex behavior: the dispersion relation also depends on the key dimensionless parameters, $St$, $\beta$, and $Re$.
@@ 4.3.1 Structural modes, $\beta=0$
Fig. 5 shows the magnitude of the dimensionless phase velocity, $|v_{p}|$, for the classical Euler–Bernoulli beam (referred to as “classical beam theory” or “CBT”) and for its nonlocal counterpart at zero fluid damping ($\beta=0$). The nonlocal attenuation visible at large $k$ is therefore due solely to the nonlocality (integral kernel); no viscous damping is present in these curves. The black curve ($\Delta=0$) reproduces the classical Euler–Bernoulli result, $|v_{p}|=k/\sqrt{St}$, and serves as the baseline. All nonlocal curves coincide with this linear behavior for long waves ($k\Delta\ll 1$), confirming that nonlocality leaves the low-frequency response unchanged. For finite horizons the curves diverge once $k\Delta=O(1)$: each nonlocal branch levels off to the analytical limit
$\lim_{k\to\infty}|v_{p}|=\frac{\pi}{\sqrt{St}\,\Delta},$ (31)
The dispersion curves in Fig. 5 are governed by two independent parameters. The Strouhal number $St$ comes from the fluid–structure coupling and fixes the reference slope: the Euler–Bernoulli limit ($\Delta=0$) gives $|v_{p}|=k/\sqrt{St}$, so increasing $St$ shifts all velocities downward in proportion to $1/\sqrt{St}$. The second parameter, the peridynamic horizon $\Delta$, controls the degree of nonlocality. For any nonzero $\Delta$ the curve departs from the classical line once $k\Delta=O(1)$ and gradually saturates to the limiting value $\pi/(\sqrt{St}\,\Delta)$. Hence, a larger horizon lowers the high-$k$ plateau, demonstrating the intrinsic high-frequency filtering of the nonlocal kernel, whereas a larger $St$ lowers the entire curve without altering its shape.
Figure 5: Dimensionless phase velocity $|v_{p}|$ versus $k/\pi$ for the classical Euler–Bernoulli beam (CBT) and the nonlocal beam with three horizon lengths $\Delta_{1}=0.02$, $\Delta_{2}=0.04$, and $\Delta_{3}=0.08$; $\beta=0$. Dashed horizontal lines show the large-$k$ asymptotes given by $\pi/\bigl{(}\sqrt{St}\,\Delta\bigr{)}$.
@@
This saturation persists for a significant range of the compliance number ($\beta<10^{3}$) and for finite Reynolds number $Re$, confirming that the elastic nonlocal kernel (not fluid damping) governs the short-wave response. When $\beta$ is small, the coupling is weak, and the dispersion becomes essentially insensitive to $Re$, provided the flow stays within the low-Reynolds-number regime relevant to microchannels, $Re\leq 1$. The break in slope at $k\Delta\approx 1$ marks the onset of a “wave-lag” effect: beyond this point, short wavelengths travel progressively slower than the prediction from Euler–Bernoulli theory. The same lag is evident in Fig. 4, where waves recorded at a fixed instant arrive later as $\Delta$ increases. Altogether, the dispersion plot illustrates three features of the fluid–peridynamic structure interaction problem: agreement with classical beam theory at long wavelengths, attenuation of high-frequency waves, and a horizon- and $St$-dependent limit on the ultimate phase velocity.
@@ 4.3.3 Damping
Next, Fig. 7 distinguishes the roles of fluid–structure coupling, peridynamic horizon, and Strouhal number in determining the spatial decay rate (damping) of harmonic waves. In the absence of coupling to the flow ($\beta=0$), the attenuation is identically zero for both the local and nonlocal beams, confirming that a purely elastic solid supports loss-free wave propagation, as expected. Introducing fluid–structure coupling ($\beta=10^{5}$) generates a finite, frequency-dependent attenuation even for the practically local case (infinitesimal horizon size).
@@
Finally, Fig. 8 isolates the influence of the Reynolds number $Re$ and the compliance number $\beta$ on the spatial damping rate of a beam with essentially local elasticity ($\Delta=10^{-5}$). Relative to the reference case ($Re=10^{-5}$, $\beta=10^{5}$), decreasing $Re$ or increasing $\beta$ reduces the attenuation in the low-frequency range but enhances it at high frequencies. In other words, a highly viscous flow or a strong fluid–structure coupling extracts little energy from slowly varying waves, but dissipates more rapidly once the oscillation rate is high enough for viscous stresses to dominate.
@@ 4.4 Damage and failure of the microchannel: Static vs dynamic loading
Figure 9: Contour plot showing the space-time evolution of the particle curvature field $C(\mathbf{x})$ as a function of dimensionless axial position $X\in[0,1]$ and dimensionless time $T\in[0,0.12]$. Colors show how $C(\mathbf{x})$ evolves over the chosen space-time interval, with warmer colors (yellow) indicating higher values and cooler colors (blue) indicating lower values. The white contour lines indicate constant $C$-values. Simulation parameters are $St=10$, $\beta=20$, $Re=0.5$, and $\Delta=1/120$.
@@
    Equivalent static loading: the beam is subjected to the static loading of a steady-state pressure distribution extracted from the same fluid–structure interaction (FSI) simulation.
@@
Fig. 10 shows that the $(St,\beta)$ plane is divided by a line in the log-log plot,
@@
Below the dividing line (i.e., for $St/\beta$ small), the ordering reverses, $\max C_{\mathrm{static}}>\max C_{\mathrm{FSI}}$, implying that the greatest curvature is attained only after the flow has settled. Hence, for all parameter pairs beneath the boundary, the curvature recorded in the steady state controls failure: the beam fails if $C_{\mathrm{cr}}<\max C_{\mathrm{static}}$ and survives otherwise.
@@
This outcome highlights how transient dynamic forces in the coupled problem can significantly amplify beam deformation, while also revealing a clear boundary between dynamic and static (potential) failure mechanisms, enabled by the proposed reduced-order fluid–peridynamic structure interaction model.
@@ 5 Conclusion
We developed a 1D model, coupling lubrication theory for a viscous flow to a nonlocal Euler–Bernoulli beam based on peridynamic theory, to simulate dynamics, up to failure, the flow-induced deformation of a channel’s compliant top wall. The key novelty of the model is to capture the nonlocal material response of the soft elastic wall using the theory of peridynamics, introducing as a parameter the dimensionless horizon $\Delta$. Additionally, a Reynolds number $Re$ quantifying flow inertia, a Strouhal number $St$ quantifying unsteady inertia of the beam, and a compliance number $\beta$ quantifying the fluid–structure coupling emerge from the model.
@@
Beyond wave dynamics, we explored the potential failure of the beam-like soft wall of the microchannel under static and dynamic conditions. We found that a boundary in the $(St,\beta)$ parameter space separates two failure-risk regimes: above this curve, transient loading is more severe than the corresponding steady load, whereas below it, failure is likely under the steady state load. The location of an “operating point” relative to this boundary, therefore, indicates whether the design conditions of microchannel operation are likely to lead to failure during its transient response to the flow or not.
Finally, while in this work we exploited the reduction to a 1D problem through depth averaging of the flow. In future work, it would be of interest to solve a 2D problem for the failure of the microchannel wall, including resolving the fracture of the wall and the subsequent dynamics of the escaping fluid. Such a simulation could be performed using the novel immersed-boundary fluid–peridynamic structure interaction solver developed by Kim et al. [54].
@@
Ziyu Wang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Visualization. Ivan C. Christov: Conceptualization, Methodology, Investigation, Writing - Original Draft, Supervision, Project administration, Funding acquisition.
@@ Appendix A Numerical scheme
In this appendix, we introduce the numerical scheme employed to solve the dynamic coupled problem of peridynamic beam deformation and fluid flow. Our approach is based on the numerical scheme for coupled problems involving classical beam deformation and flow developed in [41]. Let the time‐step size be denoted by $\Delta T$. In this approach, the compliant channel wall is modeled as a peridynamic Euler–Bernoulli beam discretized into a single array of uniformly distributed particles, each corresponding to a discrete domain volume. We assign each particle $k$ a volume $\mathcal{V}_{k}$ (with unit cross‐sectional area), mathematically equal to $\Delta X$. In this formulation, a subscript refers to the function value at the corresponding particle, whereas a superscript indicates the discrete time step index.
The peridynamic beam’s equation of motion, Eq. (16), can be discretized [25] as follows:
@@
Here, $\dot{H}$ represents the vector of vertical velocities of the “particles” along the peridynamic beam, with $\dot{(\,\cdot\,)}=\partial(\,\cdot\,)/\partial t$ and $\ddot{(\,\cdot\,)}=\partial^{2}(\,\cdot\,)/\partial t^{2}$.
@@
$\ddot{H}^{n+1}$ $=\left(St\bm{I}+\phi_{2}\Delta T^{2}\bm{K}\right)^{-1}\left\{\beta P^{n+1}\right$. (40a)
$\quad\left.-\bm{K}\left[H^{n}+\Delta T\dot{H}^{n}+\Bigl{(}\tfrac{1}{2}-\phi_{2}\Bigr{)}\Delta T^{2}\ddot{H}^{n}\right]\right\},$
@@
This scheme is unconditionally stable and exhibits second‐order accuracy when $\phi_{1}=1/2$ and $\phi_{2}=1/4$. However, choosing $\phi_{1}>1/2$ helps suppress numerically induced high-frequency oscillations [48]. In our simulations, we set $\phi_{1}=1.0$ and $\phi_{2}=0.5625$.
Meanwhile, for the flow equations, we adopt a second‐order backward difference formula in time. The dimensionless lubrication Eqs. (14) and (15) are discretized and integrated in space (using trapezoidal integration) [40, 41] to yield:
@@
$P_{k}^{n+1}$ $=\int_{1}^{X_{k}}\left[-\frac{Re}{H^{n+1}_{k}}\left(\frac{3Q_{k}^{n+1}-4Q_{k}^{n}+Q_{k}^{n-1}}{2\Delta T}\right)\right$.
$\qquad\left.+\frac{6}{5}\frac{\left(\tfrac{Q^{2}}{H}\right)^{n+1}_{k+1}-\left(\tfrac{Q^{2}}{H}\right)^{n+1}_{k-1}}{2\Delta X}-12\left(\frac{Q}{H^{3}}\right)^{n+1}_{k}\right]d\tilde{X}$. (41b)

### v2

@@ 4.3 Dispersion relation for the coupled problem
where $H_{0}$, $Q_{0}$, and $P_{0}$ denote the base state values, and $H^{\prime}$, $Q^{\prime}$, and $P^{\prime}$ represent the corresponding perturbations. Furthermore, to derive a linearized dispersion relation, we have assumed harmonic forms of the perturbations [55, 51, 9, 10] in Eqs. (22). Here, $H^{\prime}$, $P^{\prime}$, and $Q^{\prime}$ are understood as the real part of the complex-valued expressions on the right-hand sides, but we leave the real part notation understood, as is common in the literature. In these expressions, $k$ is the (dimensionless) wavenumber, and $\omega$ is the (dimensionless) frequency.
@@
where the first row arises from the PD beam equation, the second from the continuity equation, and the third row from the fluid’s axial momentum balance equation.
@@
The dispersion relation in Eq. (29) can be solved to determine the dimensionless phase velocity, $v_{p}=\Re{(\omega)}/\Re{(k)}$. Unlike the case of an uncoupled peridynamic beam or bar [55, 9, 10], where the dispersion relation depends solely on the wavenumber, the present model exhibits a more complex behavior: the dispersion relation also depends on the key dimensionless parameters, $St$, $\beta$, and $Re$.
@@ 4.3.1 Structural modes, $\beta=0$
Fig. 5 shows the magnitude of the dimensionless phase velocity, $|v_{p}|$, for the classical Euler–Bernoulli beam (referred to as “classical beam theory” or “CBT”) and for its nonlocal counterpart at zero fluid damping ($\beta=0$). The nonlocal attenuation visible at large $k$ is therefore due solely to the nonlocality (integral kernel); no viscous damping is present in these curves. The black curve ($\Delta=0$) reproduces the CBT result, $|v_{p}|=k/\sqrt{St}$, and serves as the baseline. All nonlocal curves coincide with this linear behavior for long waves ($k\Delta\ll 1$), confirming that nonlocality leaves the low-frequency response unchanged. For a finite horizon, the curves diverge once $k\Delta=O(1)$, and each nonlocal branch levels off to the limiting value
$\lim_{k\to\infty}|v_{p}|=\frac{\pi}{\sqrt{St}\,\Delta}.$ (31)
The dispersion curves in Fig. 5 are governed by two independent parameters. The Strouhal number $St$ comes from the beam inertia and fixes the reference slope: the CBT limit ($\Delta=0$) gives $|v_{p}|=k/\sqrt{St}$, so increasing $St$ shifts all velocities downward in proportion to $1/\sqrt{St}$. The second parameter, the peridynamic horizon $\Delta$, controls the degree of nonlocality. For any nonzero $\Delta$, the curve departs from the classical line once $k\Delta=O(1)$ and gradually saturates to the limiting value given in Eq. (31). Hence, a larger horizon lowers the high-$k$ plateau, demonstrating the intrinsic high-frequency filtering of the nonlocal kernel, whereas a larger $St$ lowers the entire curve without altering its shape.
Figure 5: Dimensionless phase velocity $|v_{p}|$ versus $k/\pi$ for the classical Euler–Bernoulli beam (CBT) and the nonlocal PD beam with three horizon lengths $\Delta_{1}=0.02$, $\Delta_{2}=0.04$, and $\Delta_{3}=0.08$; $\beta=0$. Dashed horizontal lines show the large-$k$ asymptotes given in Eq. (31).
@@
This saturation persists for a significant range of the compliance numbers ($\beta<10^{3}$) and for finite Reynolds number $Re$, confirming that the elastic nonlocal kernel (not fluid damping) governs the short-wave response. When $\beta$ is small, the coupling is weak, and the dispersion becomes essentially insensitive to $Re$, provided the flow stays within the low-Reynolds-number regime relevant to microchannels, $Re\leq 1$. The break in slope at $k\Delta\approx 1$ marks the onset of a “wave-lag” effect: beyond this point, short wavelengths travel progressively slower than the prediction from CBT. The same lag is evident in Fig. 4, where waves recorded at a fixed instant arrive later as $\Delta$ increases. Altogether, the dispersion plot illustrates three features of the fluid–peridynamic structure interaction problem: agreement with classical beam theory at long wavelengths, attenuation of high-frequency waves, and a horizon- and $St$-dependent limit on the ultimate phase velocity.
@@ 4.3.3 Damping
Next, Fig. 7 distinguishes the roles of fluid–structure coupling, peridynamic horizon, and beam inertia in determining the spatial decay rate (damping) of harmonic waves. In the absence of coupling to the flow ($\beta=0$), the attenuation is identically zero for both the local and nonlocal beams, confirming that a purely elastic solid supports loss-free wave propagation, as expected. Introducing fluid–structure coupling ($\beta=10^{5}$) generates a finite, frequency-dependent attenuation even for the practically local case (infinitesimal horizon size).
@@
Finally, Fig. 8 isolates the influence of the Reynolds number $Re$ and the compliance number $\beta$ on the spatial damping rate of a beam with very weak nonlocality ($\Delta=10^{-5}$). Relative to the reference case ($Re=10^{-5}$, $\beta=10^{5}$), decreasing $Re$ or increasing $\beta$ reduces the attenuation in the low-frequency range but enhances it at high frequencies. In other words, a highly viscous flow or a strong fluid–structure coupling extracts little energy from slowly varying waves, but dissipates more rapidly once the oscillation rate is high enough for viscous stresses to dominate.
@@ 4.4 Damage and failure of the microchannel: Static vs dynamic loading
Figure 9: Contour plot showing the space-time evolution of the particle curvature field $C(\mathbf{x})$ as a function of dimensionless axial position $X\in[0,1]$ and dimensionless time $T\in[0,0.12]$. Colors show how $C(\mathbf{x})$ evolves over the chosen space-time interval, with warmer colors (yellow) indicating higher values and cooler colors (blue) indicating lower values. The superimposed contours indicate constant $C$-values. Simulation parameters are $St=10$, $\beta=20$, $Re=0.5$, and $\Delta=1/120$.
@@
    Equivalent static loading: the beam is subjected to static loading based on the steady-state pressure distribution extracted from the same fluid–structure interaction (FSI) simulation.
@@
Furthermore, Fig. 10 shows that the $(St,\beta)$ plane is divided by a line in the log-log plot,
@@
Below the dividing line (i.e., for $St/\beta$ small), the ordering reverses, $\max C_{\mathrm{static}}>\max C_{\mathrm{FSI}}$, implying that the greatest curvature is attained only after the flow transients have settled. Hence, for all parameter pairs beneath the boundary, the curvature recorded in the steady state controls failure: the beam fails if $C_{\mathrm{cr}}<\max C_{\mathrm{static}}$ and survives otherwise.
@@
This outcome highlights how transient (dynamic)forces in the coupled problem can significantly amplify beam deformation, while also revealing a clear boundary between dynamic and static (potential) failure mechanisms, enabled by the proposed reduced-order fluid–peridynamic structure interaction model.
We also note that varying the peridynamic horizon $\delta$ modifies the nonlocal structural response, and thus both the quasi-static and dynamic deflection/curvature fields can vary with the dimensionless horizon $\Delta$. Consequently, even when the critical curvature threshold $C_{\mathrm{cr}}$ is held fixed, the predicted damage initiation site and onset time (i.e., the first location and time at which the computed curvature reaches $C_{\mathrm{cr}}$) may shift with $\Delta$. In the parameter regime explored here, these shifts are modest and are therefore not a primary focus of the present study.
5 Advantages and limitations of the proposed modeling approach
The present work couples a one-dimensional peridynamic Euler–Bernoulli beam model (used over the entire structural span) with a classical lubrication-theory model for the fluid. The coupling is multiphysics: the fluid enters the beam equation through a distributed transverse pressure, while the lubrication equation is closed through the gap-height kinematics implied by the beam deflection. Importantly, we do not perform any spatial decomposition of the structure into local and nonlocal subdomains (e.g., PD–FEM coupling across an internal structural interface). Consequently, well-known dynamic artifacts reported for local–nonlocal structural coupling, such as spurious wave reflections at a PD–local interface [23, 12], are not expected to arise in the present formulation.
Nonetheless, peridynamic models require careful treatment near physical boundaries due to neighborhood truncation, which can influence localized fields (and, in some cases, the static response) in the vicinity of the boundary. In the present thickness-integrated 1D peridynamic Euler–Bernoulli beam formulation, the beam ends are subject to kinematic clamped constraints (displacement and rotation) rather than traction boundary conditions. These constraints are enforced using ghost (fictitious) material points and mirror-image relations for the displacement field near each end, consistent with established peridynamic beam implementations [56, 58].
For completeness, we note that when traction-type boundary conditions must be imposed in peridynamics (typically in 2D/3D solid formulations), a common approach is the method of layers, in which the prescribed traction is represented by an equivalent body-force density over a fictitious boundary layer [25]. More recent work proposes alternative treatments that improve compatibility with local boundary-condition enforcement [19]. Such traction-imposition strategies are not required in the present 1D clamped-beam setting but may be relevant for future extensions.
In this work, we model the fluid using classical lubrication theory. This choice is based on the targeted applications of microfluidic flows, in which channels are long and slender, and easily fit into the thin-gap, low-Reynolds-number regime of interest. In these regimes and slender geometries, lubrication theory provides an accurate and computationally efficient reduced-order description of incompressible viscous flow. Our primary motivation for employing peridynamics is on the structural side, where the nonlocal peridynamic beam formulation naturally accommodates damage initiation and discontinuities, which are central to the aims of the present study. Replacing the lubrication equation with a fully peridynamic (or more broadly, nonlocal) fluid solver would substantially increase computational cost and algorithmic complexity (e.g., nonlocal enforcement of incompressibility and boundary-condition treatments), while obscuring the primary focus of this work.
Furthermore, in our simulations, we enforce the assumptions of lubrication theory through the choice of dimensionless parameters and by monitoring the maximum wall slope during time integration; if the slope exceeds a prescribed admissible threshold, the computation is terminated with a warning, indicating that the chosen parameters have entered a regime where the lubrication approximation may no longer be appropriate. Importantly, note that the deformations shown in the figures are dimensionless (scaled), which visually amplifies the apparent magnitude of deflection even when the small-slope condition is satisfied. Yet, perhaps unexpectedly, although classical lubrication is often associated with the small-slope assumptions, it has been shown to remain accurate even when channel shape changes are comparable to the channel height by comparisons to extended lubrication theory and direct numerical simulations [46].
In terms of structure, the peridynamic Euler–Bernoulli beam inherits the classical Euler–Bernoulli limitations, including the neglect of transverse shear and kinematic assumptions consistent with small rotations/deflections. The same slope-based stopping criterion is used to prevent interpretation of responses that fall outside this regime within the present model. A natural extension, if larger deformations and shear effects must be captured, is to adopt a shear-deformable beam theory (e.g., a peridynamic Timoshenko beam formulation) [57, 30].
Finally, PD and nonlocal-operator formulations for fluid flow and transport phenomena are active areas of research [11, 27, 64]. These developments motivate promising future work toward a nonlocal analogue of the Reynolds lubrication equation within the thin-film asymptotic reduction.
@@ 6 Conclusion
We developed a 1D model, coupling lubrication theory for viscous flow to a nonlocal Euler–Bernoulli beam based on peridynamic theory, to simulate the dynamics, up to failure, of the flow-induced deformation of a channel’s compliant top wall. The key novelty of the model is to capture the nonlocal material response of the soft elastic wall using the theory of peridynamics, introducing as a parameter the dimensionless horizon $\Delta$. Additionally, a Reynolds number $Re$ quantifying flow inertia, a Strouhal number $St$ quantifying unsteady inertia of the beam, and a compliance number $\beta$ quantifying the fluid–structure coupling emerge from the model.
@@
Beyond wave dynamics, we explored the potential failure of the beam-like soft wall of the microchannel under static and dynamic conditions. We found that a boundary in the $(St,\beta)$ parameter space separates two failure-risk regimes: above this curve, transient loading is more severe than the corresponding steady load, whereas below it, failure is likely under the steady-state load. The location of an “operating point” relative to this boundary, therefore, indicates whether the design conditions of microchannel operation are likely to lead to failure during its transient response to the flow or not.
Finally, in this work, we exploited the reduction to a 1D problem via depth-averaged flow. In future work, it would be of interest to solve a 2D problem for the failure of the microchannel wall, including resolving the fracture of the wall and the subsequent dynamics of the escaping fluid. Such a simulation could be performed using the novel immersed-boundary fluid–peridynamic structure interaction solver developed by Kim et al. [22].
@@
Ziyu Wang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - review & editing, Visualization. Ivan C. Christov: Conceptualization, Methodology, Investigation, Writing - Original Draft, Writing - review & editing, Supervision, Project administration, Funding acquisition.
@@ Appendix A Numerical scheme
In this appendix, we introduce the numerical scheme employed to solve the dynamic coupled problem of peridynamic beam deformation and fluid flow. Our approach is based on the numerical scheme for coupled problems involving classical beam deformation and flow developed in [54]. Let the time‐step size be denoted by $\Delta T$. In this approach, the compliant channel wall is modeled as a PD beam discretized into a single array of uniformly distributed particles, each corresponding to a discrete domain volume. We assign each particle $k$ a volume $\mathcal{V}_{k}$ (with unit cross‐sectional area), mathematically equal to $\Delta X$. In this formulation, a subscript refers to the function value at the corresponding particle, whereas a superscript indicates the discrete time step index.
The PD beam’s equation of motion, Eq. (16), can be discretized [56] as follows:
@@
Here, $\dot{H}$ represents the vector of vertical velocities of the “particles” along the PD beam, with $\dot{(\,\cdot\,)}=\partial(\,\cdot\,)/\partial t$ and $\ddot{(\,\cdot\,)}=\partial^{2}(\,\cdot\,)/\partial t^{2}$.
@@
$\ddot{H}^{n+1}$ $=\left(St\bm{I}+\phi_{2}\Delta T^{2}\bm{K}\right)^{-1}\Big\{\beta P^{n+1}$ (40a)
$\quad-\bm{K}\left[H^{n}+\Delta T\dot{H}^{n}+\Bigl(\tfrac{1}{2}-\phi_{2}\Bigr)\Delta T^{2}\ddot{H}^{n}\right]\Big\},$
@@
This scheme is unconditionally stable and exhibits second‐order accuracy when $\phi_{1}=1/2$ and $\phi_{2}=1/4$. However, choosing $\phi_{1}>1/2$ helps suppress numerically induced high-frequency oscillations [45]. In our simulations, we used $\phi_{1}=1.0$ and $\phi_{2}=0.5625$.
Meanwhile, for the flow equations, we adopt a second‐order backward difference formula in time. The dimensionless lubrication Eqs. (14) and (15) are discretized and integrated in space (using the trapezoidal rule) [20, 54] to yield:
@@
$P_{k}^{n+1}$ $=\int_{1}^{X_{k}}\Bigg[-\frac{Re}{H^{n+1}_{k}}\left(\frac{3Q_{k}^{n+1}-4Q_{k}^{n}+Q_{k}^{n-1}}{2\Delta T}\right)$
$\qquad+\frac{6}{5}\frac{\left(\tfrac{Q^{2}}{H}\right)^{n+1}_{k+1}-\left(\tfrac{Q^{2}}{H}\right)^{n+1}_{k-1}}{2\Delta X}-12\left(\frac{Q}{H^{3}}\right)^{n+1}_{k}\Bigg]d\tilde{X}$. (41b)